OpenStudy (anonymous):

PRE-CAL!!!!! HELP!!! D: okay so.. the question deals with limits. Q: lim (as x approches 0) x/x^2-x

OpenStudy (cwtan):

\[\[\lim_{x \rightarrow 0}\frac {x}{x^2-x}\]?

OpenStudy (across):

Hint: Factor the denominator and simplify.

OpenStudy (anonymous):

I know but, won't it factor to x(x-1) ? and then the lim wouldn't exist?

OpenStudy (across):

Then simplify...

OpenStudy (anonymous):

cwtan - yes you're right that's the question

OpenStudy (across):

\[\frac{x}{x(x-1)}=\frac{1}{x-1}\]

OpenStudy (anonymous):

then the limit is -1